# Point-Supervised Facial Expression Spotting with Gaussian-Based Instance-Adaptive Intensity Modeling

### v2

@@ 2.2 Point-supervised temporal action localization
Although P-FES, the focus of this study, has many similarities with P-TAL, research specifically dedicated to P-FES is extremely limited. Wang et al. [wang2025micro] investigated a similar but unique micro-expression key frame inference task. Yu et al. [yu2024weak] explored P-FES by employing a framework similar to general P-TAL methods. In this paper, instead of employing a general P-TAL framework, we propose a new two-branch framework for expression intensity modeling and class-aware apex classification.
@@ 3.1 Problem formulation
Given an untrimmed facial video $\mathbf{V}=(\mathbf{v}_{i})_{i=1}^{T}$ including a total of $T$ frames, we only annotate single timestamp for each facial expression instance and get the point set $\mathbf{Y}={(p_{i},\bm{y}_{i})}^{N}_{i=1}$, where $p_{i}$ represents the index of the annotated frame of the $i$-th expression instance during training, NN denotes the total number of ground-truth expression instances, and $\bm{y}_{i}$ denotes the multi-hot vector representing the action class (i.e., MaE and ME), respectively. Our goal is to detect as many expression instances as possible, which involves determining the boundary frames and classifying each instance as MaE or ME.
@@ 3.3 Overall framework
After obtaining the optical flow features, we input the embedded feature vectors into our proposed two-branch framework to estimate the class-agnostic expression intensity scores $\bm{a}\in\mathbb{R}^{T}$ and the class-aware apex scores $\mathbf{S}\in\mathbb{R}^{T\times C}$, where $C$ represents the number of expression classes (i.e., MaE and ME). Unlike general two-branch frameworks, we decouple the two branches and process them independently during both training and inference, rather than fusing their outputs. Note that the FES task does not involve emotion recognition; therefore, the estimated intensity scores and action scores are independent of emotional categories. In addition, the overall framework involves our proposed GIM module for soft-pseudo labeling and IAC loss for contrastive learning.
@@ 3.6.3 Inference
4 Experiments
4.1 Experimental settings
4.1.1 Datasets
4.1.2 Evaluation metrics
4.1.3 Training details
@@
4.2 Comparison with state-of-the-art methods
@@
4.3 Apex frame detection
$NMAE=\frac{1}{K}\sum_{i=1}^{K}\hat{e}_{i},$ (18)
$\hat{e}_{i}=\frac{e_{i}}{l_{i}},$ (19)
4.4 Ablation study
@@
4.4.1 Two-branch framework optimization strategy
4.4.2 Pseudo-labeling strategy
4.4.3 Handling Overlapping Gaussian Distributions
@@ 4.6 Qualitative evaluation
Intensity score results of an entire video. As shown in Fig. 6, we present the expression intensity score results for an entire video. The results indicate that our method effectively suppresses neutral noise while preserving the intensity of expressions, demonstrating its capability to distinguish between expression and neutral frames.

### v3

@@ 2.2 Point-supervised temporal action localization
Although P-FES, the focus of this study, has many similarities with P-TAL, research specifically dedicated to P-FES is extremely limited. Wang et al. [wang2025micro] investigated a similar but unique micro-expression key frame inference task. Yu et al. [yu2024weak] explored P-FES by employing a framework similar to general P-TAL methods.
In this paper, instead of employing a general P-TAL framework, we propose a new two-branch framework for expression intensity modeling and class-aware apex classification.
@@ 3.1 Problem formulation
Given an untrimmed facial video $\mathbf{V}=(\mathbf{v}_{i})_{i=1}^{T}$ including a total of $T$ frames, we only annotate single timestamp for each facial expression instance and get the point set $\mathbf{Y}={(p_{i},\bm{y}_{i})}^{N}_{i=1}$, where $p_{i}$ represents the index of the annotated frame of the $i$-th expression instance during training, $N$ denotes the total number of ground-truth expression instances, and $\bm{y}_{i}$ denotes the multi-hot vector representing the action class (i.e., MaE and ME), respectively. Our goal is to detect as many expression instances as possible, which involves determining the boundary frames and classifying each instance as MaE or ME.
@@ 3.3 Overall framework
After obtaining the optical flow features, we input the embedded feature vectors into our proposed two-branch framework to estimate the class-agnostic expression intensity scores $\bm{a}\in\mathbb{R}^{T}$ and the class-aware apex scores $\mathbf{S}\in\mathbb{R}^{T\times C}$, where $C$ represents the number of expression classes (i.e., MaE and ME). Unlike general two-branch frameworks, we decouple the two branches and process them independently during both training and inference, rather than fusing their outputs. Since the FES task does not involve emotion recognition, the estimated intensity scores and action scores are independent of specific emotional categories.
The overall framework further incorporates the proposed GIM module for soft pseudo-label generation and the IAC loss for contrastive learning. Note that the SpotFormer [deng2024spotformer] is only used as the feature extraction backbone due to its strong optical-flow-based representation learning capability for facial expression analysis. The proposed problem setting and overall framework design are not inherently tied to a specific backbone architecture. The GIM module operates on frame-level feature embeddings and assigns soft pseudo-labels based on feature distances, without imposing any constraints on how these features are extracted. Therefore, the proposed framework can be readily integrated with alternative backbone networks, including appearance-based or multimodal models, as well as other spatiotemporal models, that are capable of producing snippet-based frame-level features.
@@ 3.4 Gaussian-based instance-adaptive intensity modeling (GIM)
Although the Gaussian distribution is constructed as a symmetric intensity pattern centered at the apex frame, it serves as a structural prior rather than a rigid temporal assumption. The soft pseudo-labels are determined by the feature distance between each frame and the pseudo-apex frame instead of the temporal distance, which allows them to naturally form asymmetric intensity trajectories adaptive to each individual expression instance (analysis and visualizations are provided in Section 4.4.2 with Table V and Section 4.6 with Fig. 5). This effectively decouples expression intensity estimation from strict temporal symmetry assumptions, thereby allowing the GIM module to remain robust to asymmetric expression dynamics while preserving the interpretability of the Gaussian-based prior, that is, the apex frame corresponds to the peak intensity, and the intensity smoothly decreases on each side.
@@ 3.6.3 Inference
4 Experiments
4.1 Experimental settings
4.1.1 Datasets
4.1.2 Evaluation metrics
4.1.3 Training details
@@
4.2 Comparison with state-of-the-art methods
4.3 Apex frame detection
$NMAE=\frac{1}{K}\sum_{i=1}^{K}\hat{e}_{i},$ (18)
$\hat{e}_{i}=\frac{e_{i}}{l_{i}},$ (19)
@@ 4.3 Apex frame detection
4.4 Ablation study
4.4.1 Two-branch framework optimization strategy
4.4.2 Pseudo-labeling strategy
4.4.3 Handling Overlapping Gaussian Distributions
@@ 4.6 Qualitative evaluation
Intensity score results of an entire video. As shown in Fig. 6, we present the expression intensity score results for an entire video. The results indicate that our method effectively suppresses neutral noise while preserving the intensity of expressions, demonstrating its capability to distinguish between expression and neutral frames in most cases. However, some MEs with extremely low intensity may still be difficult to detect, as their intensity scores are weak and nearly indistinguishable from neutral motion fluctuations. These cases illustrate that distinguishing extremely subtle MEs from neutral noise remains an inherent challenge in ME spotting.